User Tools

Site Tools


handover:c1713_v1713

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
Last revision Both sides next revision
handover:c1713_v1713 [2017/12/10 17:01]
Patrick Yates
handover:c1713_v1713 [2017/12/10 23:09]
Jamie McCallum
Line 7: Line 7:
      * 00:00 UT - Experiment started OK      * 00:00 UT - Experiment started OK
      * 12:18 UT - Lost connection to drive pc, restarted both drive pc and field system pc, but no luck. Called Jamie, he is trying to fix it. Probably related to the network still not being quite right. Brett restarted some switches out at Mt Pleasant, but issues still remain. (Patrick)      * 12:18 UT - Lost connection to drive pc, restarted both drive pc and field system pc, but no luck. Called Jamie, he is trying to fix it. Probably related to the network still not being quite right. Brett restarted some switches out at Mt Pleasant, but issues still remain. (Patrick)
-     * 13:44 UT - Restarted schedule. Problem was a struggling ​dns server that was responding to queries slowly. Jamie pointed Newsmerd and Hobart to use a different DNS server which fixed the issues. Resumed with scan 344-1344. Missed scans were 344-1218a - 344-1341 (Patrick)+     * 13:44 UT - Restarted schedule. Problem was a struggling ​DNS server that was responding to queries slowly. Jamie pointed Newsmerd and Hobart to use a different DNS server which fixed the issues. Resumed with scan 344-1344. Missed scans were 344-1218a - 344-1341 (Patrick)
      * 14:51 UT - Windstow until 15:19. Had to send ''​source=stow''​ command again, as telescope was not stowing. (Patrick)      * 14:51 UT - Windstow until 15:19. Had to send ''​source=stow''​ command again, as telescope was not stowing. (Patrick)
      * 15:57 UT - Windstow until 16:59. (Patrick)      * 15:57 UT - Windstow until 16:59. (Patrick)
 +     * 17:48 UT - Windstow until 17:59. (Patrick) 
 +     * 23:06 UT - Schedule halted for cryo maintenance. (JMc)
  
 **Katherine** **Katherine**
/home/www/auscope/opswiki/data/pages/handover/c1713_v1713.txt · Last modified: 2017/12/10 23:35 by Jamie McCallum